if a person tests positive for Tuberculosis and starts treatment immediately it takes on avarage 6 months ( 183 days with a standard deviation of 14 days) of treatment until they are completely TB free .what is the probability that a person will be healthy and TB free in less than 160 days
1
Expert's answer
2020-10-15T20:01:07-0400

Let "X=" the length of treatment in days: "X\sim N(\mu, \sigma^2 )"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=183\ days, \sigma=14\ days"


"P(X<160)=P(Z<\dfrac{160-183}{14})\approx"

"\approx P(Z<-1.642857)\approx0.050206"

The probability that a person will be healthy and TB free in less than 160 days is approximately "5.0206\ \%."